OpenStudy (anonymous):

The line of symmetry for the quadratic equation y = ax2 - 8x - 3 is x = 2. What is the value of "a"?

jhonyy9 (jhonyy9):

y= a4 -16 -3 --- y =4a -19 ---

jhonyy9 (jhonyy9):

a = (19-y)/4

OpenStudy (anonymous):

answer?